# Learning Selection Cuts With Gradients

### v1

@@ Abstract
Many analyses in high-energy physics rely on selection thresholds (cuts) applied to detector, particle, or event properties. Initial cut values can often be guessed from physical intuition, but cut optimization, especially for multiple features, is commonly performed by hand, or skipped entirely in favor of multivariate algorithms like BDTs or classification networks. We revisit this problem, and develop a cut optimization approach based on gradient descent. Cut thresholds are learned as parameters of a network with a simple architecture, and can be tuned to achieve a target signal efficiency through the use of custom loss functions. Contractive terms in the loss can be used to ensure a smooth evolution of cuts as functions of efficiency, particle kinematics, or event features. The method is used to classify events in a search for Supersymmetry, and the performance is compared with traditional classification networks. An implementation of this approach is available in a public code repository and python package.
@@ 1 Introduction
Rule-based data analysis is a common starting point for many studies in high-energy physics. Specific physics processes often motivate phase space requirements, such as cuts on particle momenta or event activity, that are based on physical principles. Optimization of those cut values, using significance or other metrics, can lead to event selections that are easily understood in the context of the physics signature under study. While an analysis based only on one-dimensional cuts will rarely match the performance of a multivariate optimization that fully exploits correlations between particle or event features, the transparency of cut-based procedures, and the relatively straightforward interpretations of the phase space captured by the cuts, make them attractive as a starting point for data analysis.
@@
Cut-based event selections are also useful in other ways. A set of one-dimensional cuts constructs a contiguous volume in feature space that is naturally robust against training samples that do not fully span all features, while deep neural networks often identify islands in feature space whose definitions can be more sensitive to the details of the training procedure. This can be especially undesirable in cases where the training data are derived from samples of simulated events that may not perfectly model the event features used during evaluation. The signal efficiency of a single volume in feature space may be easier to calibrate in such cases, which can be useful in measurements where excellent control over signal efficiency is required. Cut-based selections are also commonly used in event generators and hardware triggers, where more complicated selections based on neural networks are challenging to implement (though these challenges are quickly being overcome, see [Jia:2024ysq] for a recent review of machine learning in hardware).
@@
Despite the utility of cut-based approaches, tools to derive optimal cut values are limited, and mostly consist either of TMVA’s kCuts routine [hoecker2009tmvatoolkitmultivariate], or brute-force scans of feature space (see e.g. the optimization package). The kCuts algorithm is a common starting point for cut-based analyses, and derives cuts that maximize the background rejection at different signal efficiency working points. Our work was partially motivated by the observation that kCuts tunes the cuts for each target efficiency separately, such that small changes in the target efficiency can result in large changes in cuts. Subsets of events with similar (but not identical) features can also lead to very different cut values when treated separately in training. We desired an algorithm that was relatively insensitive to the exact value of the target signal efficiency, and that could provide well-behaved (ideally monotonically-varying) cuts as functions of common event properties like pileup or particle momenta. This also presented an opportunity to implement such an algorithm using modern machine learning libraries.
@@
This work presents an approach to binary classification where optimized cuts are extracted from the parameters of a trained neural network with an extremely simple architecture. Training is performed using gradient descent, requiring the modeling of cuts using differentiable functions. The application of a set of selection criteria is modeled as a product of sigmoid-activated linear transformations of input features, where the biases and weights of the linear transforms can be translated into cuts. We refer to this as the "cuts as biases in networks" (CABIN) approach. Custom loss functions are used to derive cuts for specific target efficiencies, and to regulate the evolution of cut values across efficiency targets and other event properties. We test the performance of this technique using a randomly generated dataset, and study the stability of the training procedure for different network hyperparameters. We also investigate the performance of this method in a study of LHC sensitivity to SUSY models using samples of simulated events. Ref. [Watts:2024wes] studies this problem in the context of cuts used as preselection requirements prior to the use of a neural network classifier, and confronts some of the challenges we discuss here, such as the expression of cuts as differentiable functions. Our work focuses on the optimization of cuts, without including any additional multivariate classifiers, and introduces contractive terms to control the behavior of cuts across working points.
@@ 2 Learning cuts as biases
We consider the problem of assigning binary labels to events $\mathbf{x_{i}}$ of a set of events. Each $\mathbf{x_{i}}$ can be represented as a vector of input features: $\mathbf{x_{i}}=(x_{i,1},...,x_{i,k},...,x_{i,m})$ where we use $k$ to index the features from $1$ to $m$, and $i$ to index the events.
In a typical neural network trained for binary classification, the $m$ features are treated as inputs for a fully-connected feed-forward network that has a single output, representing a sum of linear transformations of the input values passed through one or more non-linear activation functions, that provides a score for a given event $\mathbf{x_{i}}$. After training, events from one category will cluster at one end of the network output score distribution, while events from the other category cluster at the other end of the range of scores. A threshold on the output score distribution can then be chosen based on signal efficiency, background rejection, significance, or other metrics. In this work we are interested in deriving thresholds on the input features, not on the summary score.
@@
If the weights are chosen to be unity (up to a sign) then the bias $b_{k}$ represents the cut value for feature $k$. As a parameter of a linear layer in a model, the optimal value of the cut can be learned through training.
@@ 2.1 Differentiable cuts
The problem with learning a set of rectangular cuts, which is equivalent to a product of Heaviside step functions $\Theta_{k}(x)$, is that step functions cannot be used for gradient-based training. We address this by approximating $\Theta_{k}(x)$ with the logistic function, $\sigma(x)$, where the inputs are multiplied by some factor $n$:
@@
For finite $n>1$, this function is differentiable and provides a good approximation of the Heaviside function as long as $x$ is not near 00, the threshold used to distinguish between categories. Figure 1 shows the behavior of the logistic function for different values of $n$.
@@
The score goes to zero if the event fails any cut, and only converges to a value near one if the event passes all cuts, and by a sufficient margin that they are away from the sigmoid turn-on for every input feature.
The final output of this algorithm is a score, not a binary decision. The scores asymptotically approach 00 or $1$ for events that fail or pass the selection cuts, respectively, and for sufficiently large values of $n$ the sum of scores for events that are evaluated by the network closely matches the number of events that pass the cuts learned in the training procedure. On the other hand, values of $n$ that are too large can make learning unstable, as gradients vanish quickly when $x^{\prime}$ values move away from zero.
@@ 2.2 Loss Functions
The usual loss function used for two-category classification is binary cross entropy (BCE), which penalizes incorrect classification of any objects in the dataset. Another approach, used by the kCuts algorithm, is to tune the cuts to yield a specified target efficiency, and then maximize the background rejection for that efficiency. In order to reproduce the TMVA procedure, which is a common starting point for many HEP analyses, we implement a loss function, referred to here as the "efficiency loss", based on deviations from a target signal efficiency, in addition to penalizing non-zero background efficiency:
@@
where $\varepsilon_{\mathrm{signal}}^{\mathrm{estimated}}$ is the sum of output scores for true signal events, normalized to the number of true signal events, and $\varepsilon_{\mathrm{background}}^{\mathrm{estimated}}$ is the sum of output scores for true background events, normalized the number of true background events. Positive-valued hyperparameters $\alpha$ and $\beta$ can be used to control the relative importance of signal efficiency vs background rejection during training. As discussed in Section 2.2, the choice of $n$ in equation 4 influences how closely $\varepsilon_{\mathrm{signal}}^{\mathrm{estimated}}$ approximates the actual signal efficiency obtained by applying the cuts derived from the network parameters. Since $\varepsilon_{\mathrm{background}}^{\mathrm{estimated}}\geq 0$, the loss defined in equation 5 is convex.
Implementation of a custom efficiency loss function also allows the introduction of contractive terms that regulate the behavior of cuts individually and across different optimization regions. The first contractive term we consider is based on the magnitude of the cuts themselves:
@@
where $b_{k}$ are the biases, $m$ is the number of features, and $\gamma$ is a positive-valued hyperparameter that controls the influence of this loss term during training. This contractive term provides a penalty based on the magnitude of the cut that avoids having cut values that are far from the training distributions.
@@
3 Efficiency scan networks
The one-to-one linear networks described above, in combination with the custom loss function that trains a network to yield a target signal efficiency, allows the construction of a collection of simultaneously-trained networks that span multiple efficiency targets. An associated loss function, called the efficiency loss, incorporates efficiency loss terms for each of the sub-networks, as well as contractive terms that regulate the behavior of the cut values across different sub-networks. The contractive terms are implemented as penalties for cuts that are not between cuts for neighboring target efficiencies:
@@ 4 Implementation
We implement the CABIN method in the cabin python package [cabin] using pytorch [pytorch]. A simple, one-to-one linear network is constructed in which $m$ input features feed to $m$ nodes in a single hidden layer, with each input connected to one and only one hidden node. The network is illustrated in Fig. 2. The input feature first undergoes the linear transformation in equation 1, and then goes through a logistic activation function. The outputs of the hidden layer are multiplied to produce a single output score for each event:
@@
We find that $n=15$ results in output scores for signal events that are close enough to unity that the sum of the output scores is a good estimate of the number of signal events that pass cuts derived from the weights and biases of the network. The weights can either be fixed by the user, or learned as part of the training procedure. We prefer setups in which the weights $w_{k}$ are fixed to ±$\pm 1$, such that they only encode whether the cut represents a "greater than" or "less than" requirement.
@@
Figure 2: A one-to-one linear network used to learn cuts as biases. The version illustrated here has fixed weights, which determine whether the associated bias corresponds to a "greater-than" or "less-than" cut.
The training procedure is sensitive to the initial values assigned to the weights and biases. If the weights and biases result in cuts that are far from the bulk of the distribution then the gradients of the loss function can vanish, leading to poor training convergence. Initializing cut values in the middle of the distribution helps with this. We use the StandardScaler function in scikit-learn [scikit-learn] to rescale the inputs to have zero mean and unit width. The biases are then initialized to 0, which helps avoid vanishing gradients at the start of training.
For comparison with more traditional approaches, we also implement a simple version of a linear network, with $m$ inputs, one output, and no hidden layers, illustrated in Figure 3. The network output is calculated as:
$\mathrm{score}_{i}=\sigma\left(\sum_{k=0}^{m-1}w_{k}x_{i,k}\right).$ (9)
This network, with trainable weights and no biases, has the same number of free parameters ($m$) as the one-to-one linear network in equation 8 that has trainable biases and fixed weights. A BCE loss function is used to evaluate this model.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$y=\sigma\left(\sum_{k}w_{k}x_{k}\right)$$w_{1}$$w_{2}$$w_{3}$$w_{4}$
Figure 3: A simple linear network, with trainable weights and no bias term, used for comparison with the one-to-one linear network that learns cuts as biases.
Networks are trained with the torch.optim.SGD optimizer with a learning rate of 0.5, processing data in a single batch for 100 to 200 epochs. Since these networks are simple by construction, training times are typically fast, and no attempts were made to optimize the training hyperparameters or tune for fast convergence of the loss. The hyperparameters included in the loss functions are summarized in Table 1, along with values used during training.
@@
5 Data
We use a dataset of 20,000 events with ten features, where event properties are generated by randomly selecting values from Gaussian distributions with randomly-chosen means and widths. No correlations between features are implemented. Feature distributions for this dataset, referred to here as the Gaussian dataset, are shown in Fig. 4. We use 80% of the Gaussian dataset for training, and reserve 20% for testing.
Figure 4: Feature distributions in the Gaussian dataset for signal (red) and background (blue) events used for initial development of the CABIN approach. Features are Gaussian-distributed with assorted means and widths. No correlations between features are implemented.
A second dataset is used to study the behavior of the CABIN approach in a collider physics context. We consider the problem of discriminating events arising from Supersymmetric interactions from those of the Standard Model (SM). In particular, we simulate the production of pairs of sleptons, scalar SUSY partners of SM leptons, which are separated in mass from the lightest neutralino, here also the lightest SUSY partner, by approximately 20 GeV. The final state includes two opposite-sign leptons from the slepton decays, along with missing transverse energy ($E_{\mathrm{T}}^{\mathrm{miss}}$) from the invisible neutralinos. In such scenarios, it is often advantageous to require a hard jet, nominally from initial state radiation, which boosts the SUSY system. The SM production of $WW$ events, with both bosons decaying leptonically, provides an irreducible background that is difficult to suppress. Example Feynman diagrams of the SUSY signal and irreducible SM background process is shown in Fig. 5.
We generate events for signal and background processes using MadGraph5_aMCNLO version 5.2.7.3 [Alwall:2014hca] using the NNPDF 2.3 leading order parton distribution functions [Ball:2012cx]. Events are showered with Pythia version 8.244 [Sjostrand:2014zea], and passed through a parameterized simulation of the ATLAS detector response implemented in Delphes version 3.5.0 [deFavereau:2013fsa, Selvaggi:2014mya, Mertens:2015kba]. Events in the SUSY search are pre-selected to have two leptons, at least one energetic jet ($p_{\mathrm{T}}>100$ GeV), and substantial missing transverse energy ($E_{\mathrm{T}}^{\mathrm{miss}}>300$ GeV). After preselection, we consider the following event features:
  • $E_{\mathrm{T}}^{\mathrm{miss}}$
  • the $p_{\mathrm{T}}$ of the leading jet
  • the transverse mass of each lepton and the $E_{\mathrm{T}}^{\mathrm{miss}}$:
  • the separation $\Delta R$ of the leptons from each other:
Figure 6: Distributions of kinematic features used to classify SUSY and SM $WW$ events.

### v2

@@ Abstract
Many analyses in high-energy physics rely on selection thresholds (cuts) applied to detector, particle, or event properties. Initial cut values can often be guessed from physical intuition, but cut optimization, especially for multiple features, is commonly performed by hand, or skipped entirely in favor of multivariate algorithms like BDTs or neural networks. We revisit this problem, and develop a cut optimization approach based on gradient descent. Cut thresholds are learned as parameters of a network with a simple architecture, and can be tuned to achieve a target signal efficiency through the use of custom loss functions. Contractive terms in the loss can be used to ensure a smooth evolution of cuts as functions of efficiency, particle kinematics, or event features. The method is used to classify events in a search for Supersymmetry, and the performance is compared with common classification tools. An implementation of this approach is available in a public code repository and python package.
@@ 1 Introduction
Rule-based data analysis is a common starting point for many studies in high-energy physics. Specific physics processes often motivate phase space requirements, such as cuts on particle momenta or event activity, that are based on physical principles. Optimization of those cut values, using significance or other metrics, can lead to event selections that are easily understood in the context of the physics signature under study. While an analysis based only on one-dimensional cuts will rarely match the performance of a multivariate optimization that fully exploits correlations between particle or event features, the transparency of cut-based procedures, and the relatively straightforward interpretations of the phase space captured by the cuts, make them attractive as a starting point for data analysis. Cuts have been used by many recent ATLAS and CMS searches for supersymmetry to select events for primary signal regions [1, 2, 3, 4, 5, 6]. Another use case for cuts is to define "discovery" regions that enable future reinterpretation [7], often by complementing a primary selection based on boosted decision tree (BDT) or neural network outputs.
@@
Cut-based event selections are also useful in other ways. A set of one-dimensional cuts constructs a contiguous volume in feature space that is naturally robust against training samples that do not fully span all features, while deep neural networks identify islands in feature space whose definitions can be more sensitive to the details of the training procedure. This can be especially undesirable in cases when training data are derived from samples of simulated events that may not perfectly model event features used for evaluation. The signal efficiency of a single volume in feature space may be easier to calibrate, which can be useful in measurements where precise knowledge of the signal efficiency is required. For example, the ATLAS experiment uses cuts to discriminate photons from non-prompt backgrounds [8, 9]. Cut-based selections are also commonly used in event generators and hardware triggers, where more complicated selections based on neural networks are challenging to implement (though these challenges are quickly being overcome, see [10] for a recent review of machine learning in hardware).
@@
Despite the utility of cut-based approaches, tools to derive optimal cut values are limited, and mostly consist either of TMVA’s kCuts routine [11], or brute-force scans of feature space (see e.g. the optimization package). TMVA’s kCuts algorithm is a common starting point for LHC physicists optimizing cut-based analyses; it derives cuts that maximize the background rejection at different signal efficiency working points. Our work was partially motivated by the observation that kCuts tunes the cuts for each target efficiency separately, such that small changes in the target efficiency can result in large changes in cuts. Subsets of training events with similar (but not identical) features can also lead to very different cut values. We desired an algorithm that was relatively insensitive to the exact value of the target signal efficiency, and that could provide well-behaved (ideally monotonically-varying) cuts as functions of common event properties like pileup or particle momenta. This also provided an opportunity to implement such an algorithm using modern machine learning libraries.
@@
This work presents an approach to binary classification where optimized cuts are extracted from the parameters of a trained neural network with an extremely simple architecture. Training is performed using gradient descent, requiring selection cuts to be modeled with differentiable functions. The application of a set of selection criteria is modeled as a product of sigmoid-activated linear transformations of input features, where the biases and weights of the linear transforms can be translated into cuts. We refer to this as the "cuts as biases in networks" (CABIN) approach. Custom loss functions are used to derive cuts for specific target efficiencies, and to regulate the evolution of cut values across efficiency targets and other event properties. We test the performance of this technique using a dataset of simulated LHC events, and study the stability of the training procedure for different network hyperparameters.
Ref. [12] studies this problem in the context of cuts used as preselection requirements prior to the use of a neural network classifier, and confronts some of the challenges we discuss here, such as the expression of cuts as differentiable functions. Our work focuses on the optimization of cuts, without including any additional multivariate classifiers, and introduces contractive terms to control the behavior of cuts across working points. However, both Ref. [12] and this work develop methods that are fully differentiable, and can be included as components of larger analysis pipelines that make use of auto-differentiation for training and inference.
@@ 2 Learning cuts as biases
We consider the problem of assigning binary labels to records $\mathbf{x_{i}}$ in a data sample, here assumed to be particle collision events. Each $\mathbf{x_{i}}$ can be represented as a vector of input features: $\mathbf{x_{i}}=(x_{i,1},...,x_{i,k},...,x_{i,m})$ where we use $k$ to index the features from $1$ to $m$, and $i$ to index the events.
In a typical neural network trained for binary classification, the $m$ features are treated as inputs for a fully-connected feed-forward network that has a single output score $y$, representing a sum of linear transformations of the input values passed through one or more non-linear activation functions, that provides a score for a given event $\mathbf{x_{i}}$. After training, events from one category will cluster at one end of the network output score distribution, while events from the other category cluster at the other end of the range of scores. A threshold on the output score distribution can then be chosen based on signal efficiency, background rejection, significance, or other metrics. The goal of this work is to derive thresholds on the input features, instead of a single threshold on the output score.
@@
If the weights are chosen to be unity (up to a sign) then the bias $b_{k}$ represents the cut value for feature $k$. As a parameter of a linear layer in a differentiable model, the optimal value of the cut can be learned through training.
@@ 2.1 Differentiable cuts
The problem with learning a set of rectangular cuts, expressed as a product of Heaviside step functions $\Theta_{k}(x)$, is that step functions cannot be used for gradient-based training. We address this by approximating $\Theta_{k}(x)$ with the logistic function, $\sigma(x)$, where the inputs are multiplied by some factor $n$:
@@
This function is differentiable, and for finite $n>1$ provides a good approximation of the Heaviside function as long as $x$ is not near 0, the threshold used to distinguish between categories. Figure 1 shows the behavior of the logistic function for different values of $n$.
@@
The score goes to zero if the event fails any cut, and only converges to a value near one if the event passes all cuts (and by a sufficient margin that they are away from the sigmoid turn-on for every input feature).
The final output of this algorithm is a score, not a binary decision. The score asymptotically approaches 0 or $1$ for events that fail or pass the selection cuts, respectively, and for sufficiently large values of $n$ the sum of scores for events that are evaluated by the network closely matches the number of events that pass the cuts learned in the training procedure, motivating a large value of $n$. On the other hand, values of $n$ that are too large can make learning unstable, as gradients vanish quickly when $x^{\prime}$ values move away from the bulk of the training data.
@@ 2.2 Loss Functions
The usual loss function used for two-category classification is binary cross entropy (BCE), which penalizes incorrect classification of any objects in the dataset. Another approach, used by TMVA’s kCuts algorithm, is to tune the cuts to yield a specified target efficiency, and then maximize the background rejection for that efficiency. In order to reproduce the kCuts procedure, which is a common starting point for many HEP analyses, we implement a loss function, referred to here as the "efficiency loss", based on deviations from a target signal efficiency ($\varepsilon_{\mathrm{signal}}^{\mathrm{target}}$), in addition to penalizing non-zero background efficiency:
@@
where $\varepsilon_{\mathrm{signal}}^{\mathrm{estimated}}$ is the sum of output scores for true signal events, normalized to the number of true signal events, and $\varepsilon_{\mathrm{background}}^{\mathrm{estimated}}$ is the sum of output scores for true background events, normalized to the number of true background events. Positive-valued hyperparameters $\alpha$ and $\beta$ can be used to control the relative importance of signal efficiency vs background rejection during training. As discussed in Section 2.2, the choice of $n$ in Equation 4 influences how closely $\varepsilon_{\mathrm{signal}}^{\mathrm{estimated}}$ approximates the actual signal efficiency obtained by applying the cuts derived from the network parameters. Since $\varepsilon_{\mathrm{background}}^{\mathrm{estimated}}\geq 0$, and the term scaled by $\alpha$ is also positive, the loss defined in Equation 5 is convex, and will have a finite global minimum.
Implementation of a custom efficiency loss function also allows the introduction of contractive terms [13] that regulate the behavior of cuts individually and across different optimization regions. The first contractive term we consider is based on the magnitude of the cuts themselves:
@@
where $b_{k}$ are the biases, $m$ is the number of features, and $\gamma$ is a positive-valued hyperparameter that controls the influence of this loss term during training. This contractive term provides a penalty based on the magnitude of the cut, and prevents learning cut values that are far from the training distributions.
@@
3 Data
We assess the CABIN approach using simulated collider physics data, focusing on the problem of discriminating events arising from Supersymmetric interactions from those of the Standard Model (SM). In particular, we simulate the production of pairs of sleptons, scalar SUSY partners of SM leptons, which are separated in mass from the lightest neutralino, here also the lightest SUSY partner, by approximately 20 GeV. The final state includes two opposite-sign leptons from the slepton decays, along with missing transverse energy ($E_{\mathrm{T}}^{\mathrm{miss}}$) from the invisible neutralinos. In such scenarios, it is often advantageous to require a hard jet, nominally from initial state radiation, which boosts the SUSY system. The SM production of $WW$ events, with both bosons decaying leptonically, provides an irreducible background that is difficult to suppress. Example Feynman diagrams of the SUSY signal and irreducible SM background process is shown in Fig. 2. The ATLAS collaboration recently studied this final state in Ref. [7], where it used cut-based selections to test for generic excesses as well as BDT-based selections to constrain specific Supersymmetric scenarios.
We generate events for signal and background processes using the mapyde toolkit [14], with MadGraph5_aMCNLO version 5.2.7.3 [15] and NNPDF 2.3 leading order parton distribution functions [16]. Events are showered with Pythia version 8.244 [17], and passed through a parameterized simulation of the ATLAS detector response implemented in Delphes version 3.5.0 [18, 19, 20]. Events in the SUSY search are pre-selected to have two leptons with $p_{\mathrm{T}}>10$ GeV, at least one energetic jet ($p_{\mathrm{T}}>100$ GeV), and substantial missing transverse energy ($E_{\mathrm{T}}^{\mathrm{miss}}>300$ GeV). The generated samples have 88k signal events and 33k background events after preselection, 80% of which are used for training, and 20% of which are used for testing.
Training procedures for all methods use the following event features:
  • $E_{\mathrm{T}}^{\mathrm{miss}}$
  • the $p_{\mathrm{T}}$ of the leading jet
  • the transverse mass of each lepton and the $E_{\mathrm{T}}^{\mathrm{miss}}$:
  • the separation $\Delta R$ of the leptons from each other:
Figure 3: Distributions of kinematic features used to classify SUSY and SM WWWW events.
@@ 4 Implementation
The CABIN method is tested using the SUSY dataset described above, and compared with three alternative approaches: a cut-based selection tuned using TMVA’s kCuts algorithm; a simple, fully-connected neural network; and a BDT.
4.1 CABIN method
We implement the CABIN method in the cabin python package [21] using pytorch [22]. The CABIN method optimizes a sequence of orthongonal 1-dimensional cuts to partition the $m$-dimensional feature space into signal-enriched and background-enriched regions.
A one-to-one linear network is constructed in which $m$ features ($m=7$ in the SUSY dataset) are inputs to $m$ nodes in a single hidden layer, with each input connected to one and only one hidden node. The network is illustrated in Fig. 4. The input feature first undergoes the linear transformation in equation 1, followed by a logistic activation function. The outputs of the hidden layer are multiplied to produce a single output score for each event:
@@ 4.1 CABIN method
We find that $n=15$ results in output scores for signal events that are close enough to unity that the sum of the output scores is a good estimate of the number of signal events that pass the derived cuts, while still avoiding vanishing gradients during optimization. The weights $w_{k}$ can either be fixed by the user, or learned as part of the training procedure. We prefer setups in which the weights are fixed to $\pm 1$, such that they only encode whether a cut represents a "greater than" or "less than" requirement.
@@
Figure 4: A one-to-one linear network used to biases ($b_{k}$) that correspond to cuts. The sign of the weight ($w_{k}$) determines whether the bias term corresponds to a "greater-than" or "less-than" cut; for simplicity we usually fix them to $\pm 1$.
The training procedure is sensitive to the initial values assigned to the weights and biases. If the initial weights and biases result in cuts that are far from the bulk of the distribution then the gradients of the loss function can vanish, leading to poor training convergence. Initializing cut values in the middle of the distribution helps with this. We use the StandardScaler function in scikit-learn [23] to rescale the inputs to have zero mean and unit width. The biases are then initialized to 0, which helps avoid vanishing gradients at the start of training.
CABIN networks are trained with the torch.optim.SGD optimizer with a learning rate of 0.5, processing data in a single batch for 100 to 200 epochs. Since these networks are simple by construction, training times are typically fast, and no attempts were made to optimize the training hyperparameters or tune for fast convergence of the loss. The hyperparameters included in the loss functions are summarized in Table 1, along with values used during training. CABIN networks are studied with both a pure-BCE loss function as well as a loss that targets a specific signal efficiency.
@@
4.1.1 Efficiency scan networks
The one-to-one linear networks described above, in combination with the custom loss function that trains a network to yield a target signal efficiency, allows the construction of a collection of simultaneously-trained networks that span multiple efficiency targets. The loss function for this collection of networks incorporates efficiency loss terms for each of the sub-networks, as well as contractive terms that regulate the behavior of the cut values across different sub-networks. The contractive smoothing terms are implemented as penalties for cuts that are not between cuts for neighboring target efficiencies:
The efficiency scan networks are implemented as a collection of one-to-one linear networks correlated through the $\varepsilon$ term in Table 1. The efficiency scan networks implemented here have target efficiencies ranging from 10% to 90% in steps of 10%, as well as targets at 1%, 5%, 93%, and 95% to explore the behavior at more extreme values of the target signal efficiency.
4.2 Simple Neural Network
For comparison, we also implement a simple version of a neural network, with $m$ inputs, one output, and no hidden layers, illustrated in Figure 5. In this approach the $m$-dimensional feature space is partitioned with a hyperplane with optimized orientation in m1m-1 dimensions. The network output is calculated as:
$\mathrm{score}_{i}=\sigma\left(\sum_{k=0}^{m-1}w_{k}x_{i,k}\right).$ (9)
This network, with trainable weights and no biases, has the same number of free parameters ($m$) as the one-to-one linear network in equation 7 that has trainable biases and fixed weights. A BCE loss function is used with this model. The learning rate was set to 0.1, compared to 0.5 in the CABIN networks, to avoid oscillations in the loss.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$y=\sigma\left(\sum_{k}w_{k}x_{k}\right)$$w_{1}$$w_{2}$$w_{3}$$w_{4}$
Figure 5: A simple linear network, with trainable weights and no bias terms.
4.3 TMVA kCuts
The TMVA kCuts method is also used to train cuts for the example datasets. Signal and background events have unit weights in the training. The kCuts routine can be configured to use MC sampling, a Genetic Algorithm, or Simulated Annealing. The Genetic Algorithm is described in the manual [11] as having the best performance, and is used here. The training time for kCuts varies linearly with the product of the population size, number of steps, and number of cycles (roughly corresponding to the product of batches and epochs when training a neural network). We tested two configurations:
  1. 1.
    population = 200; steps = 10; cycles = 5
  2. 2.
    population = 500; steps = 20; cycles = 6
The behaviors of the output cuts and ROC curves of the two approaches are qualitatively similar. We use configuration #2 for comparison with the CABIN approach. We also use the EffSel method to target specific efficiencies, the VarProp=FSmart option to categorize cuts, and employ the CreateMVAPdfs option to rescale the data for processing internally.
4.4 Boosted Decision Tree
Finally, a BDT is trained to provide a ceiling for the performance of cut-based approaches. We use the xgBoost [24] toolkit for training a BDT to separate signal from background. BDT’s with a binary:logistic objective function and a logloss metric were trained with a range of estimators (500 through 2000, in steps of 500) and depths (2 through 7, in steps of 1) and tested for convergence, accuracy within a test sample, and overtraining through comparisons of training and test losses. Signs of overtraining were observed for depths beyond 3, and overtraining was exacerbated for large numbers of estimators. A configuration of 500 estimators and a depth of 3 was used for comparisons with the CABIN approach.